### v1

# Modular Techniques for Synthetic Long-Context Data Generation in Language Model Training and Evaluation

Abstract
The ability of large language models (LLMs) to process and reason over long textual inputs is critical for a wide range of real-world applications. However, progress in this area is significantly constrained by the absence of high-quality, diverse, and verifiable long-context datasets suitable for both training and evaluation. This work introduces a modular, extensible framework for synthetic long-context data generation via prompt-based interaction with LLMs. The framework supports multiple training and alignment objectives, including Supervised Fine-Tuning (SFT), Direct Preference Optimization (DPO), and Group Relative Policy Optimization (GRPO). It encompasses four core generation paradigms: multi-turn conversational dialogues, document-grounded input-output pairs, verifiable instruction-response tasks, and long-context reasoning examples. Through templated prompting, a model-agnostic architecture, and metadata-enriched outputs, the proposed approach facilitates scalable, controllable, and purpose-aligned dataset creation for advancing long-context capabilities in LLMs.
1 Introduction
Large Language Models (LLMs) have achieved impressive capabilities across a broad spectrum of natural language processing (NLP) tasks [7]. However, many real-world applications—such as legal analysis, scientific summarization, and multi-turn user assistance—require effective processing of long textual contexts that significantly exceed the input lengths observed during standard model training [5, 14]. Despite architectural advances such as extended-context transformer variants, the performance and reliability of LLMs in long-context settings remain fundamentally constrained by the scarcity of high-quality, task-aligned datasets.
Manual construction of such datasets is labor-intensive, costly, and frequently lacks the diversity and granularity required to support advanced training objectives such as reward modeling, preference learning, and chain-of-thought alignment. Furthermore, the absence of verifiability—i.e., the ability to evaluate whether outputs can be grounded in the provided input context—reduces the applicability of many existing datasets for alignment strategies like Direct Preference Optimization (DPO) or Group Relative Policy Optimization (GRPO) [6].
This work introduces a modular and extensible framework for generating synthetic long-context data using LLM-based prompting pipelines. Our approach supports multiple post-training objectives, including SFT, DPO, GRPO, Reward Modeling, evaluation and other tasks, by systematically targeting four long-context generation modalities: (i) multi-turn chat simulation, (ii) document-grounded task construction, (iii) verifiable instruction-response pairs, and (iv) long-context reasoning tasks.
Figure 1: Overview of the long-context data generation framework. Prompt templates and metadata are used to guide four task-specific generators, producing outputs suited for training and evaluation.
The broader role of data quality in LLM effectiveness is illustrated in Figure2. Here, the synthetic data generation process feeds directly into fine-tuning and evaluation stages, creating a feedback loop that amplifies the importance of precise, diverse, and verifiable long-context examples. Poor data quality at this stage can propagate downstream, degrading alignment, factuality, and reasoning performance.
Figure 2: The LLM lifecycle highlighting the centrality of data generation in downstream training quality. Our framework targets the data generation sub-process (in blue), ensuring that synthetic examples are diverse, task-relevant, and verifiable.
To guide our investigation, we pose the following research questions:
  • RQ1: How can synthetic data generation be systematically adapted to create high-quality, long-context datasets aligned with diverse training objectives (e.g., SFT, DPO, GRPO, evaluation)?
  • RQ2: What design decisions in prompting, metadata conditioning, and task formulation impact the controllability, coherence, and verifiability of long-context outputs?
  • RQ3: Can a unified generation framework scale across multiple domains and reasoning tasks while maintaining fidelity and generalization?
By answering these questions, we aim to support the broader shift toward data-centric LLM development, where the quality and structure of long-context data become primary levers for improving model alignment and evaluation.
2 Related Work
Long-context language modeling has become a central challenge as applications increasingly require processing and reasoning over extended sequences. Early benchmarks such as LongBench [12] and Loong [15] have systematically evaluated LLMs’ abilities to handle long-context understanding, multi-document question answering, and chain-of-reasoning tasks. However, these benchmarks often rely on a mix of real and synthetic data, with the synthetic portion primarily used to fill gaps where natural long-context data is sparse [16].
Name Year Task Description
WildLong [5] 2025 SFT, DPO, GRPO Framework for scalable, realistic synthetic long-context instruction data generation, supporting multi-turn, document-grounded, and verifiable tasks.
LongPO [17] 2025 DPO Self-evolution framework for LLMs using short-to-long preference optimization and synthetic long-context data.
AgoraBench [18] 2024 Eval Evaluation suite for assessing LLMs as synthetic data generators, focusing on diversity, quality, and downstream utility of generated long-context data.
LongAlign [6] 2024 DPO, GRPO Recipe for aligning LLMs to long-context tasks using preference optimization and Group Relative Policy Optimization with synthetic data.
LONGATTN [14] 2024 Eval Method for selecting long-context training data via token-level attention, improving efficiency and performance of LLMs on extended inputs.
Loong [15] 2024 Eval Multi-document QA benchmark for evaluating LLMs’ long-context and cross-document reasoning abilities.
S3Eval [13] 2024 Eval Synthetic, systematic evaluation suite for LLMs, focusing on long-context reasoning and instruction-following capabilities.
LongBench [12] 2023 Eval Comprehensive benchmark for long-context understanding and generation, covering QA, summarization, and retrieval tasks across various domains.
LongSkywork [16] 2023 SFT Training recipe for efficiently extending the context length of language models using synthetic and real data.
Table 1: Summary of Key Long-Context Papers
Synthetic data generation for LLM training and evaluation has emerged as a practical solution to the scarcity of high-quality, diverse long-context datasets [13, 8, 16]. Recent work has demonstrated that prompt-driven pipelines can create instruction-tuned datasets for fine-tuning and alignment, supporting tasks such as summarization, multi-turn dialogue, and compositional reasoning. Methods like SynAlign [9] and SoftSRV [10] further improve the distributional match and diversity of synthetic data through distribution matching and soft prompt optimization, respectively.
Alignment and reward modeling approaches—including RLHF [1], Direct Preference Optimization (DPO) [3], and Group Relative Policy Optimization (GRPO) [11]—increasingly depend on high-quality, verifiable long-context data. However, most existing pipelines rely on web-scraped or loosely structured sources, which may lack the task alignment and verifiability required for robust reward modeling. Recent advances, such as LongAlign and LongPO [17], propose recipes for self-evolving LLMs using preference optimization and synthetic data generation, demonstrating that carefully curated synthetic datasets can significantly enhance long-context alignment.
Evaluation of synthetic data generation frameworks has also matured. For example, AgoraBench [18] provides standardized settings and metrics to compare LLMs’ data generation abilities, revealing that model selection and output format have significant impacts on downstream effectiveness. Studies show that synthetic data can improve both training and evaluation, provided it is diverse, high-quality, and task-aligned.
Despite these advances, most prior work either focuses on benchmarking, static corpus construction, or generic instruction tuning. Our approach distinguishes itself by introducing a modular, objective-aware framework for synthetic long-context data generation, emphasizing verifiability, task-specific prompt engineering, and extensibility across alignment paradigms.
3 Methodology
The data generation framework introduced in this work is modular and extensible, designed to support long-context training and evaluation for multiple LLM alignment objectives including SFT, DPO, etc.
Figure 3: End-to-end pipeline for long-context data generation.
The overall pipeline is summarized in Figure3, which illustrates the process of synthetic long-context data construction and validation. The pipeline begins with a specified use case, which may be instantiated either as a high-level scenario or as a more complex instruction. These elements serve as the foundation for generating realistic task settings. In the scenario branch, the process proceeds to a dedicated long-context data generation module, which constructs extended textual inputs designed to emulate real-world long-form contexts. In parallel, the complex instruction branch augments the long-context data with additional constraints, and this is appended with long context data for a meaningful input to the llm.
The resulting long-context input serves as the central artifact for subsequent stages. To ensure task fidelity, the input undergoes an initial rule-based validation step, where criteria such as satisfying the desired token length or other user-specified constraints are checked. Inputs that pass these conditions are forwarded to an LLM-based answer generation component, which produces candidate responses conditioned on the validated context. These candidate responses are stored in a response buffer, which is then subjected to further evaluation. At this stage, an LLM-based judge assesses the quality, coherence, and faithfulness of the responses with respect to the original input. Finally, the validated outputs are collected and stored as final data, completing the pipeline.
This design integrates deterministic checks (e.g., token length thresholds) with LLM-based evaluators, thereby balancing precision with flexibility. By combining scenario-driven context generation, instruction-based augmentation, automated response synthesis, and layered evaluation, the pipeline enables scalable production of high-quality long-context datasets suitable for alignment, training, and benchmarking.
3.1 Multi-Turn Chat Generation
This module produces long-form, multi-turn conversations between a user and one or more AI assistants (e.g., Assistant 1, Assistant 2, …, Assistant $n$). The objective is to synthesize extended dialogue contexts that preserve semantic coherence, instructional relevance, and dynamic turn-taking across speakers. The full data generation flow is illustrated in Figure4.
Figure 4: Recursive multi-turn conversation generation. A scenario is sampled and initialized to produce an initial conversation seed. The dialogue is expanded by recursively feeding the output back as seed input, producing progressively longer sequences across multiple assistant turns.
Step 1: Scenario Repository and Sampling. The pipeline begins with a curated Scenario Database. For each datapoint, a scenario $s$ is sampled uniformly (or by a task-aware prior) to ensure topical and stylistic diversity. The sample fixes the domain, roles, and high-level objective for the conversation instance.
Step 2: Random Scenario Initialization. The scenario $s$ is instantiated into a concrete setting by randomly drawing entity placeholders, constraints, and seeds (e.g., names, tools, subgoals, time/place). This yields an initialized prompt $x_{0}$ specifying the user goal and any guardrails (format, knowledge scope).
Step 3: Multi-User Conversation Generator. Conversation participants are configured as one User role and $n\geq 1$ Assistant roles (default $n=1$). A turn function $g(\cdot)$, given the dialogue history $H_{t}$ and role $r\in\{\text{User},\text{Assistant}_{1},\ldots,\text{Assistant}_{n}\}$, produces the next message $m_{t+1}=g(H_{t},r)$. Refer to the sample prompt used in the appendix section Prompt Templates and Structure.
Step 4: Inner Loop- Build $K$ Turns per Segment. A segment of $K$ alternating user–assistant turns is generated:
  1. 1.
    User issues a goal-conditioned message based on $x_{0}$ and history $H_{t}$.
  2. 2.
    Assistant $j$ replies (optionally rotating among assistants if $n>1$).
  3. 3.
    The new message is appended to the conversation buffer: $H_{t+1}=H_{t}\oplus m_{t+1}$.
This process repeats until the segment reaches $K$ turns or a target token budget.
Step 5: Length Control Within Segment. After each turn, the cumulative token length $L(H_{t})$ is computed. If the segment has not reached the desired length, additional turns or expanded replies are generated. This corresponds to the “repeat $K$ times” loop in the figure.
Step 6: Outer Loop- Stitch $N$ Segments. Steps 4–5 are repeated for $N$ segments, with optional variations in subgoals, tools, or assistant personas. The segments are concatenated to yield an $N\times K$-turn long conversation.
Step 7: Conversation Store. The accumulated turns are stored in a Conversation Store, along with metadata:
$\{\text{scenario\_id},n,N,K,\text{seeds},\text{token counts}\}.$
This supports filtering, replay, or regeneration.
Step 8: Rule-Based Validation. Deterministic checks are applied prior to finalization:
  • Length: Total tokens fall within the target range; per-turn maxima are respected.
  • Structure: Strict alternation of roles with exactly $N\times K$ turns; no empty messages.
  • Format: Required prefixes/suffixes are present; citations/placeholders are resolved.
  • Policy: Optional filters for PII redaction, domain constraints, and toxicity limits.
Conversations failing any rule are repaired or regenerated.
Step 9: Final Long-Context Data. Validated conversations are exported as Final Long-Context Data, each item consisting of the full transcript plus metadata for downstream use (training, preference optimization, or evaluation).
Notation Recap.
  • $n$: number of assistants (default $n=1$).
  • $K$: number of turns per segment (inner loop).
  • NN: number of segments concatenated (outer loop).
Thus, the total conversation length is approximately $N\times K$ turns, further adjusted via token-level validation.
3.2 Document-Grounded Task Generation
The document-grounded generation pipeline comprises a sequence of methodologically structured components that produce semantically dense, long-context input–output pairs. These components jointly control instruction complexity and contextual depth, which are essential for training and evaluating models under long-context regimes. Below we describe each stage.
1. Structured, Multi-Field Prompt Construction. Each instance is assembled from multiple fields using prompt templates (e.g., Listing 5), including:
  • a business scenario (e.g., insurance claim analysis);
  • task-specific text-generation guidance and a brief rationale;
  • a geographically grounded context (e.g., France, Brazil);
  • auxiliary metadata such as verifiability requirements, user tone, and behavioral tags.
This configuration yields a multi-dimensional prompt that simulates real-world task diversity and ensures contextual richness. Expanding across these dimensions increases both token count and semantic density prior to generation.
2. Recursive Scenario Enrichment for Complexity. To simulate sophisticated reasoning or planning tasks, the system optionally applies a secondary template (e.g., Listing 6) that transforms a simple prompt into a multi-layered scenario. The enriched prompt:
  • introduces multiple agents, stakeholders, or roles;
  • specifies competing objectives and constraints (regulatory, financial, technical);
  • embeds edge cases and technical subtasks into the narrative.
This recursive transformation substantially increases prompt length and the depth of reasoning required for downstream outputs.
3. Long-Context Document Generation. Given the enriched scenario, the generate_user_content template (Listing 10) synthesizes a document-scale passage (the “source document”) that serves as ground truth context. The template enforces:
  1. 1.
    Locale and language constraints: the document is set outside the U.S. and must be in English only;
  2. 2.
    Controlled variation: synonym substitutions, sentence restructuring, and phrase reordering to promote lexical and structural diversity;
  3. 3.
    Length control: a target token budget $L_{\text{target}}$; if $L<L_{\text{target}}$, the system requests expansion (adding sections, examples, or appendices) until $L_{\text{min}}\leq L\leq L_{\text{max}}$;
  4. 4.
    Safety and policy checks: automatic regeneration on violations (non-English tokens, disallowed locales).
The result is a semantically rich, self-contained document that grounds subsequent tasks.
4. Document-Grounded Instruction Generation. From the source document, the create_instructions template (Listing 11) induces a complex, verifiable instruction. The instruction:
  • explicitly references the document as the sole source of truth;
  • imposes multiple constraints (e.g., word limits, perspective switching, formatting requirements);
  • requires a structured JSON output, enabling automatic validation.
We optionally canonicalize the instruction by normalizing key names, enforcing unambiguous units, and banning vague phrases (e.g., “about,” “roughly”) to improve determinism.
5. Document-Grounded Response Generation & Compliance Checking. Conditioned on both the document and the instruction, the create_response template (Listing 12) produces a model response that must strictly satisfy all constraints. We then apply rule-based validators:
  1. 1.
    Format validation: JSON parseability; presence and types of required keys; schema conformance if a schema is provided;
  2. 2.
    Factual grounding: string/number exact-match checks for document entities, amounts, dates, and identifiers referenced in the response;
  3. 3.
    Constraint satisfaction: word-count bounds, perspective requirements, and ordering/formatting rules;
  4. 4.
    Length and policy filters: response tokens within limits; language/locale constraints satisfied.
Failures trigger automatic repair or regeneration with targeted feedback (e.g., “missing key summary_assistant” or “user section not exactly 50 words”). The final artifact for each instance is the triplet
$(\text{document},\ \text{instruction},\ \text{response})$
augmented with metadata (scenario, country, tone, validator logs, token counts).
3.3 Verifiable Instruction-Response Generation
The generation pipeline for verifiable instruction-response data closely mirrors the structure of document-grounded prompting, including the use of multi-field templates and complexity enrichment. However, it introduces additional steps aimed at ensuring response verifiability and contextual grounding, both critical for alignment methods such as Group Relative Policy Optimization (GRPO).
  • Verifiability-Aware Prompt Construction: In addition to standard instruction and scenario metadata, prompts explicitly incorporate a factual or semi-factual long context. These are crafted such that the correctness of a model’s output can be explicitly determined from the input alone.
  • Correctness Variation: For each prompt, one or more completions are generated by varying model temperature or conditioning. Some responses are intentionally structured to diverge from the ground truth, enabling preference comparisons or binary reward labeling.
  • Response Annotation: Each response is labeled with a verifiability indicator (e.g., fully_verifiable, partially_verifiable, non_verifiable). This enables use in reward modeling setups where traceable correctness is a key learning signal.
3.3.1 Verifiable Rewards via JSON Schema Matching:
To support automatic evaluation of factual consistency and structural adherence, each prompt-response pair can optionally include a machine-readable JSON schema. This schema specifies required fields, expected data types, and additional constraints such as token length. Reward assignment is determined by comparing model outputs against the schema along the following dimensions:
  • Field-Level Match: The output must include all required fields with content that logically or exactly aligns with reference values.
  • Type Compliance: Each field’s value must match the specified data type, such as strings, integers, booleans, or date formats.
  • Length Constraints: Certain fields may require minimum or bounded token counts (e.g., summary fields must contain at least 50 tokens).
This structured validation enables transparent, programmatic scoring of generated outputs and supports integration with reward modeling strategies such as binary classification, preference ranking, or score-based filtering in GRPO and DPO pipelines.
This structured validation allows for deterministic and interpretable scoring, aligning closely with reward modeling and binary classification objectives in GRPO-style pipelines.
These additions render the framework suitable for supervised or preference-based training regimes requiring factual consistency, field-specific structure, and answer traceability.
3.4 Long-Context Reasoning Generation
Building upon the prompt-rich design of document-grounded tasks, the long-context reasoning generator emphasizes multi-step inferencing across extended inputs. The foundational methodology—multi-field templating, complexity injection, and structured packaging—remains unchanged, but is extended to support explicit reasoning supervision.
  • Reasoning Task Injection: Prompts are constructed to elicit reasoning across multiple spans within the long input context. These may involve causal chains, comparisons, or compositional queries. Optional hints (e.g., "think step-by-step") are embedded to encourage chain-of-thought outputs.
  • Intermediate Trace Generation: The model is prompted to provide not just final answers, but also the reasoning trace. This enables training with structured reasoning supervision, and supports evaluation under stepwise consistency metrics.
  • Logical Postprocessing: Outputs are optionally passed through filtering mechanisms that validate coherence, stepwise validity, or redundancy. Inconsistent or hallucinated reasoning steps are flagged for exclusion.
These extensions make the pipeline effective for training and evaluating models on reasoning-intensive tasks, particularly within the DPO or chain-of-thought (CoT) alignment paradigms.
3.5 Nuance and Diversity Generation
In addition to structural and task-specific variations, our framework incorporates targeted strategies for enhancing realism and diversity across generated conversations and prompts. This includes dynamic identity generation and role-based interaction types that contribute to lexical, semantic, and cultural richness in the data.
Name Personalization Based on Geography To introduce cultural and geographic specificity, all named entities—such as users, agents, or personas—are dynamically generated using the Faker library. Names are conditioned on the country metadata selected for the scenario (e.g., faker.name() with locale fr_FR for French users or en_IN for Indian agents). This introduces subtle lexical and syntactic variations in:
  • Salutations and honorifics (e.g., “Monsieur Dupont” vs. “Mr. Patel”)
  • Named entity placement and coreference
  • Conversation tone and personalization cues
Role Diversity Through Chat Type Variations Three primary configurations are supported to simulate different interaction patterns:
  1. 1.
    User $\leftrightarrow$ Assistant (LLM) — the canonical model-instruction exchange, where the assistant responds to user queries or tasks.
  2. 2.
    User $\leftrightarrow$ User (Person) — simulates human-to-human conversations such as support escalation, peer troubleshooting, or social interactions.
  3. 3.
    Hybrid (Mixed Roles) — alternating persona roles where LLM assistants adopt different tones or expertise levels, or where multiple human-style agents interact with a shared goal.
Controlled Diversity Injection To maintain balance and realism, role types are sampled probabilistically, and each conversation’s metadata includes:
  • Role configuration (e.g., user_assistant, peer_chat)
  • Persona tags for each participant (e.g., agent_type=legal, user_mood=anxious)
  • Geographic context and language hints
These techniques together enrich the training corpus with global, heterogeneous, and sociolinguistically varied samples—beneficial for both model generalization and safety alignment objectives.
4 Limitations and Considerations
While the framework offers a powerful toolkit for long-context generation and alignment, several limitations remain:
  • Model Dependency: Output quality is contingent on the capability of the base LLM. Weaker models may produce incoherent or factually incorrect samples.
  • Synthetic Distribution Bias: The diversity of generation depends on prompt variability. Without careful template design, outputs may exhibit style or structure redundancy.
  • Domain Generality: Most templates are tuned for general-purpose domains; domain-specific adaptation (e.g., legal, medical) requires targeted prompt engineering.
  • Evaluation Costs: Verifying long-context reasoning and dialogue naturalness still demands human oversight in many cases, limiting scalability of full evaluation.
Appendix: Implementation Notes
Prompt Templates and Structure
Prompt templates are constructed using string interpolation and conditional logic, with slots filled from scenario metadata (e.g., tone, task type, instruction behavior). Templates are designed for maximal clarity and consistency, with some randomization in phrasing to enhance output diversity. All templates are stored in a version-controlled schema to ensure reproducibility.
[…]
Text Generation Guidance: {{text_generation_guidance}}
Text Generation Guidance Explanation: {{text_generation_guidance_explanation}}
Country: {{country}}
[…]
If the input is:
[…]
- Text Generation Guidance: Case task generation
- Country: Brazil
[…]
"Create a case task related to automated refund processing for a retail company in São Paulo, Brazil, ensuring that different refund request categories are handled efficiently."
[…]
To ensure diversity and prevent repetition, introduce **controlled randomness** while maintaining coherence and accuracy. Apply **at least two** of the following transformations in each response:
1. **Synonym Substitutions** Replace at least **five key words** with appropriate synonyms while preserving meaning.
2. **Sentence Restructuring** Modify the structure of at least **two sentences** while keeping intent intact.
3. **Reordering Phrases** Slightly alter the order of key phrases without changing the scenarios meaning.
4. **Mild Redundancy** Introduce an occasional **extra descriptive phrase** or clarification to add variation.
[…]
- Ensure the generated scenario is contextually relevant, detailed, and realistic.
- The scenario must take place outside the U.S. If the given country is missing or ambiguous, randomly select a non-U.S. country.
- The output **must be in English only**.
- **Do not include any non-English words, phrases, characters, or scripts.**
- **Chinese, Japanese, or any other non-English language elements are strictly prohibited.**
- If any non-English words appear, regenerate the response ensuring complete adherence to this rule.
[…]
Figure 5: Sample prompt for Creating A Scenario
[…]
The transformed scenario must be set in a country and city outside the United States.
[…]
**Simple scenario:** Generate a case task related to Automated refund processing for a company in Hyderabad, India with various sections.
**Transformed complex scenario:** Develop a case task related to automated refund processing by integrating:
- **Regulatory compliance** (e.g., local financial laws and global data protection policies).
- **Multiple payment gateways** with differing transaction rules.
- **Fraud detection mechanisms** to prevent misuse and false claims.
- **Multi-tier customer dispute resolution** involving legal, financial, and technical teams.
- **System scalability considerations** to handle high transaction volumes efficiently.
[…]
#### **Country:** {{country}}
[…]
- Expand the scenario by incorporating **at least four** additional layers of complexity, such as **technical, regulatory, financial, operational, security, and user experience challenges**.
- Ensure the transformation introduces **multiple stakeholders** (e.g., compliance teams, technical teams, financial auditors, legal advisors, customer support).
- **Avoid generic expansions**—each added challenge should be highly specific and tailored to the context.
- Maintain realism by ensuring **the complexity aligns logically** with the nature of the original scenario.
- **Do not include phrases like "Complex Scenario" or "Transformed Scenario"** in the output.
[…]
To ensure diversity and prevent repetition, introduce **controlled randomness** while maintaining coherence and accuracy. Apply **at least two** of the following transformations in each response:
1. **Synonym Substitutions** Replace at least **five key words** with appropriate synonyms while preserving meaning.
2. **Sentence Restructuring** Modify the structure of at least **two sentences** while keeping intent intact.
3. **Reordering Phrases** Slightly alter the order of key phrases without changing the scenarios meaning.
4. **Mild Redundancy** Introduce an occasional **extra descriptive phrase** or clarification to add variation.
[…]
- The output **must be in English only**.
- **Do not include any non-English words, phrases, characters, or scripts.**
- **Strictly prohibit** cross-lingual elements, particularly **Chinese or Japanese characters**.
- If any non-English elements appear, **automatically regenerate** the response while ensuring full compliance.
- Dont use U.S cities like Austin, Texas, Denver in the response.
[…]
Figure 6: Sample prompt for Creating A Scenario Complex
You are a skilled AI assistant. Generate a natural, detailed, and realistic conversation between two or three participants based on the given scenario in a specific country.
[…]
1. The user first interacts with assistant-1 (L1 support).
2. After a few exchanges, assistant-1 hands off the conversation to assistant-2 (L2 support).
3. When the handoff happens:
* The user may re-explain the issue OR
* Assistant-2 may continue seamlessly, assuming they have chat history.
4. The issue may either be resolved or remain unsolved after assistant-2’s response.
[…]
The first name provided is always the user.
The second name provided is always the assistant-1.
The third name provided is always the assistant-2.
[…]
1. Country : {{country}}
2. Scenario: {{conversation_scenario}}
3. User Tone: {{user_tone}} (may include unorganized information, confused, abusive words, spelling mistakes, informal language, and noise such as emails, URLs, or irrelevant text)
4. Assistant Tone: Always formal, polite, and patient. The assistant should ask only necessary and precise questions when seeking information.
5. Assistant-2 awareness of chat : {{chat_awareness}} (True: Assistant-2 has full context; False - The user must re-explain the issue)
6. Assistant-2 solved the issue :{{solution_status}} (True - Assistant-2 provides a convincing solution; False - The issue remains unresolved.)
[…]
1. The dialogue should feel completely natural and human-likeno one should suspect its machine-generated.
2. Use realistic phrasing, contractions, and informal structures where appropriate.
3. The user should not sound roboticthey may hesitate, backtrack, or provide unnecessary details.
4. The user can use abusive or threatening words to force the assistant to get what he/she wanted.
5. The assistant should be professional but sound human, not overly scripted.
6. Add minor pauses, filler words (e.g., "um," "you know"), and corrections where necessary for authenticity.
7. Ensure the flow of conversation makes senseresponses should be logical and adaptive.
8. **Each generated conversation must have at least one unique element** (e.g., misunderstanding, humor, unexpected turn).
9. **Vary the users style across generations**—sometimes clear, sometimes disorganized, sometimes emotional.
[…]
"{{user_name}}": "<user conversation>",
"{{assistant_1_name}}": "<assistant-1 conversation>"
"{{user_name}}": "<user conversation>",
"{{assistant_1_name}}": "<assistant-1 conversation>"
"{{user_name}}": "<user conversation>",
"{{assistant_2_name}}": "<assistant-2 conversation>"
"{{user_name}}": "<user conversation>",
"{{assistant_2_name}}": "<assistant-2 conversation>"
[…]
Output must be in English. No cross-lingual languages particularly Chinese or Japanese are not allowed.
Generate a realistic scenario set in a French city. Include
Figure 7: Sample prompt for Creating Conversation
You are an advanced AI specializing in generating structured and complex instructions. Based on the given conversation, create a detailed and specific instruction that requires deep analysis of the conversation.
[…]
2. The task should be logically complex, requiring the AI to process information in a structured manner.
3. The instruction should not be generic; it must demand deep analysis and precise formatting.
4. The instruction must explicitly ask for the output in a proper JSON format.
[…]
2. It may include word limits, format constraints, or multiple perspectives when applicable.
3. Ensure the instruction challenges the AI to produce a nuanced response.
4. The instruction should not be generic; it should require deep analysis or structured output.
5. Output only the generated instruction. Do not include any explanation, metadata, or additional text.
[…]
Figure 8: Sample prompt for Creating Conversation Instruction
[…]
Instruction : {{instructions}}
[…]
2. Ensure that every condition, constraint, and formatting rule mentioned in the instructions is fully met.
3. If the instructions specify a particular format (e.g., JSON, XML, bullet points, etc.), the output must strictly follow it.
4. Maintain accuracy, coherence, and completeness in the generated response.
5. Do not omit or alter any part of the instructionsensure 1006. Generate only the final responsedo not include explanations, processing notes, or metadata.
7. If the user conversation contain abusive or inappropriate words, do not use them in the response.
7. Output must be in English. No cross-lingual languages particularly Chinese or Japanese are not allowed.
[…]
Figure 9: Sample prompt for Creating Conversation Response
[…]
#### **Country:** {{country}}
[…]
- Ensure the generated scenario is **contextually relevant, detailed, and realistic**.
- The output **must be in English only**.
- **Do not include any non-English words, phrases, characters, or scripts.**
- **Chinese, Japanese, or any other non-English language elements are strictly prohibited.**
- If any non-English words appear, **regenerate** the response to ensure full adherence.
- Do not use any U.S. cities (e.g., Austin, Texas; Denver, Colorado). If the provided country is missing or unclear, randomly select a non-U.S. country.
[…]
To ensure diversity and prevent repetition, introduce **controlled randomness** while maintaining coherence and accuracy. Apply **at least two** of the following transformations in each response:
1. **Synonym Substitutions** Replace at least **five key words** with appropriate synonyms while preserving meaning.
2. **Sentence Restructuring** Modify the structure of at least **two sentences** while keeping intent intact.
3. **Reordering Phrases** Slightly alter the order of key phrases without changing the scenarios meaning.
4. **Mild Redundancy** Introduce an occasional **extra descriptive phrase** or clarification to add variation.
[…]
Each time, apply one of the following subtle stylistic variations:
- A slightly **formal tone**
- A **conversational** and engaging tone
- A **descriptive style** with sensory details
- A **neutral, straightforward** approach
[…]
Dont use U.S cities like Austin, Texas, Denver in the response.
[…]
Figure 10: Sample prompt for Creating Long context Document
[…]
Based on the given text, create a detailed and specific instruction that requires deep analysis of the conversation.
[…]
2. The task should be logically complex, requiring the AI to process information in a structured manner.
3. The instruction should not be generic; it must demand deep analysis and precise formatting.
4. The instruction must explicitly ask for the output in a proper JSON format.
[…]
2. It can include word limits, format constraints, or multiple perspectives when applicable.
3. Ensure the instruction challenges the AI to produce a nuanced response.
4. The instruction should not be generic; it should require deep analysis or structured output.
5. Output only the generated instruction. Do not include any explanation, metadata, or additional text.
[…]
Figure 11: Sample prompt for Creating Long context Document Instruction
You are an advanced AI capable of processing complex instructions with high accuracy.
Given a conversation and a set of instructions, generate a response that strictly adheres to every detail of the instructions without any omissions.
[…]
Instruction : {{instructions}}
[…]
2. Ensure that every condition, constraint, and formatting rule mentioned in the instructions is fully met.
3. If the instructions specify a particular format (e.g., JSON, XML, bullet points, etc.), the output must strictly follow it.
4. Maintain accuracy, coherence, and completeness in the generated response.
5. Do not omit or alter any part of the instructionsensure 1006. Generate only the final responsedo not include explanations, processing notes, or metadata.
7. Output must be in English.
8. No cross-lingual languages particularly Chinese or Japanese are not allowed.
[…]
Figure 12: Sample prompt for Creating Long context Document Response